OpenStudy (anonymous):

Find the value of y. -2y+16+6y=44

OpenStudy (anonymous):

-2y+16+6y=44= 4y+16=44= 4y=28= y=7

OpenStudy (anonymous):

6 - 2 = 4, 44 - 16 = 28 28/4 = 7